OpenStudy (anonymous):

Differentiate y=x^3. Hence find to points on y=x^3+6 where the tangent has gradient 75....

OpenStudy (anonymous):

two* :):)

hartnn (hartnn):

so could you differentiate y=x^3 ?

OpenStudy (anonymous):

yes 3x^2

hartnn (hartnn):

gradient of tangent at a point is the derivative of the curve at that point so, 75 = 3x^2 will give you 2 values of x co-ordinate where the gradient of tangent is 75

OpenStudy (anonymous):

So then i just put the x points back into the original equation to get the y points????????

hartnn (hartnn):

find 2 values of x first! then plug them in y=x^3+6 to get 2 y co-ordinates hence u get 2 points

OpenStudy (anonymous):

Sorry.....struggling to find the two x point....................

hartnn (hartnn):

75 =3x^2 x^2 =25 x = \(\pm ???\)

OpenStudy (anonymous):

+or - 5.............sorry made a dumb mistake

hartnn (hartnn):

no problem :) plug in x=+5 in y=x^3+6 find y then plug in x=-5 in y=x^3+6 find y and you get your 2 points :)

OpenStudy (anonymous):

thanks heaps again...........life saver...........(-5,-119), (5,131)

hartnn (hartnn):

those are correct :) most welcome ^_^